Imagine the market demand and supply for chicken is such that Qd = 2,000 - 300P and Qs = 800 + 200P. To aid chicken producers, the U.S. government agrees to put a price support on chickens of $3.00 per unit. If this price support goes into effect, how many chickens will the government be forced to buy? How much will the government spend on this policy?
Answer #1

Qd = 2000- 300P

Qs =800+ 200P

When government agrees to put a price support on chickens of $3 per unit. Then ,

Qd = 2000-300(3)= 1100 chickens.

Qs= 800+ 200(3)= 1400 chickens.

Government will be forced to buy (1400-1100)= 300 chickens.

The government will spend (300)(3)= $900 on this policy.